the number of positive integers whose square is a factor of 2000 is out of
A) 3
B) 6
C)10
D) 12
E)20
Thank you for your help!
I really don't have a clue how to do this, neither do my friends. And we all don't know WHAT intergers are. Thanks!!

Answer by Edwin McCravy(20056)

Positive integers are the counting numbers, the numbers we use to
count with: 1,2,3,4,5,6, etc. (ad infinitum)

1 is a factor of every positive integer, and the square of 1 is 1,
and so 1 fits the requirements.  So one such integer is

1. 1

To find the others,

we break 2000 down into prime factors.

We get 24×53

The perfect squares we can make of those are 

2.  22, 
3.  24 = (22)2 = 42 
4.  52

And when we multiply two squares, we get another square, so we can
also get:
  
5. 2252 = (2×5)2 = 102,
6. 4252 = (4×5)2 = 202,

So there are 6 such integers:

1.  1 is such because 12 = 1 = 2000÷2000
2.  2 is such because 22 = 4 = 2000÷500
3.  4 is such because 42 = 16 = 2000÷125
4.  5 is such because 52 = 25 = 2000÷80
5.  10 is such because 102 = 100 = 2000÷20
6.  20 is such because 202 = 400 = 2000÷5

Edwin